Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 54,000 $ 197,500
May 44,000 $ 172,000
June 64,000 $ 223,000
July 74,000 $ 248,500

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 44,000 machine-hour level of activity is:

Utilities (variable) $ 61,600
Supervisory salaries (fixed) 44,000
Maintenance (mixed) 66,400
Total overhead cost $ 172,000

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $248,500 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $248,500 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 49,000 machine-hours?

  • Required 1

1. Estimate how much of the $248,500 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $248,500 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) (Do not round intermediate calculations.)

Maintenance cost in July $130,500
  • Required 2

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Y = + X
  • Required 3

Express the company’s total overhead cost in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Variable Cost per Machine-Hour Fixed Cost
Utilities cost
Supervisory salaries cost
Maintenance cost
Total overhead cost $0.00 $0
Y = + X
  • Required 4

What total overhead cost would you expect to be incurred at an activity level of 49,000 machine-hours? (Do not round intermediate calculations.)

Total overhead costs
Answer #1

1)Answer is 100,900

Level of Activity
44000 MH 74000 MH
   Total factory overhead cost      172,000.00    248,500.00
Deduct:
Utilities cost @ $1.40 per MH      61,600.00    103,600.00
Supervisory salaries        44,000.00      44,000.00
Maintenance cost        66,400.00    100,900.00

2)

Machine
Hours
Maintenance cost  
High activity level                   74,000.00                    100,900.00
Low activity level                 44,000.00                       66,400.00
Change                 30,000.00                       34,500.00

Variable cost per unit = 34,500/30,000 = 1.15

Fixed cost = 66,400 - (44,000 x 1.15) = 15,800

Cost Formula = Y = $15,800 + 1.15X

3)

Variable Rate per Machine-Hour   Fixed Cost
Maintenance cost                              1.15                       15,800.00
Utilities cost                            1.40
Supervisory salaries                         44,000.00
Totals                            2.55                       59,800.00

Cost Formula =Y = 59,800 + 2.55 X

4)

Fixed cost      59,800.00
Variable costs: $2.55 per MH × 49,000 MHs      124,950.00
Total overhead costs      184,750.00